# Large Language Models, scientific knowledge and factuality: A systematic analysis in antibiotic discovery

### v2

@@ 5.1 LLMs’ biases towards compounds and fungi
Bias towards Aspergillus - the most cited fungus in PubMed.
@@
We observe that the models are biased towards certain compounds (RQ1). For instance, BioGPT generates the fungus Aspergillus as the answer in most of the prompts (79%, 45/57), which is rarely correct (only 11 times; 19%) (Table S.5). We accredit this bias to the imbalance in the training corpus, as the Aspergillus is a ubiquitous fungus in home and hospital environments. The genus Aspergillus, consists of a few hundred opportunistic species found in various climatic conditions, of which in humans, A. fumigatus is the most common and life-threatening airborne pathogen, which is particularly important among immunocompromised hosts [40]. The PubMed search outputs almost 60k references related to Aspergillus. The most cited compound from our analysis Ergosterol outputs  6000, second Aphidicolin  2500. Thus, we argue Aspergillus is over-represented in the training corpus and occurs as a statistically most probable fungus to be related to chemicals.
@@
Rare chemical compound confused with an overrepresented biological process.
@@
Another example, Conidiogenone was consistently not recognised by the best models (GPT-4, ChatGPT, BioGPT-large, Galactica). We argue that the models confuse Conidiogenone with conidiogenesis or conidiation (RQ1). Conidiogenesis is a widespread morphogenetic process that filamentous fungi undertake for dispersion. Conidiation is also important for pathogenicity of phytopathogens. Understanding the cellular mechanism of conidiation is a highly relevant research topic, thus it has been extensively studied. Conidiogenone is associated with nine results in a PubMed search, conidiogenesis with 346 and conidiation with 6047. The strongest and widespread stimulus for conidiation among filamentous fungi is the exposure of hyphae to the air. Often the generated texts focused on this aspect of the process, suggesting that the model encoded the signal for the representation of the process instead of the chemical compound.
Output not focused on fungi despite the prompt context.
@@
In the case of Ergosterol, which is a compound widely known for its significance in the scientific community, the LLM specifically trained for the bio-domain (BioGPT-Large) outperforms the rest (RQ3). It generated seven factual Ergosterol-fungus relations from 11 prompts (Table S.7). For the other models, the relation to fungus was either minimally addressed or not considered at all, particularly for the models that have generally demonstrated superior efficacy (i.e., GPT-4 with 0/11 and ChatGPT with 4/11 factual responses). We observe that GPT-4 and ChatGPT focus on Ergosterol’s biosynthesis in fungi and relevance as an antifungal target, rather than on the producing fungi, despite the prompt that explicitly defines such context. Importantly, these two models do not generate text as a continuation of the prompt, but they generate new paragraphs.
@@ 5.2 Limited ability of LLMs as Knowledge Bases
In this study, we provide evidence of the limitations in the determination of factual knowledge from LLMs in the text generation tasks. Best models were able to provide only 6/10 of the factual description of chemical entities, and only 5/10 of the factual description of chemical-fungus relations. Despite small sample size in our evaluation, it clearly shows that the performance is below the threshold allowed for a systematic and reliable application (RQ1, RQ4).
@@ 5.3 High dependency and variability due to prompt design choices
A desirable property of a LLM is to a consistent answer regardless of the prompts design. For instance, the prompt ‘{entity 1} is produced by fungi, such as’ should be equivalent to ‘{entity 1} is isolated from fungi, such as’ because it refers to the same biological relation. However, we observe a variation in the output as the evidence of sensitivity for prompt design (RQ1). Similarly to Masked Language Models [41], in order to fully exploit GPTs capabilities, prompt engineering is required. In our study, we observe that providing context in the prompt improves the performance. Of note, ChatGPT and GPT-4 are the least sensitive and produce new paragraphs instead of finishing the prompt (RQ3).
@@ 5.4 Llama 2, despite large size, responses lack factuality and specificity
GPT-4 outperforms ChatGPT (as well the rest tested models) in terms of factuality of the answer (RQ3). It scores 70% (7/10) and 43.6% (48/110) for chemical entity recognition and chemical-fungi relation recognition, respectively. In general, the factual knowledge of the entire generated text was higher (still significantly higher in GPT-4), but the output did not contain the relation that we asked for, so they were considered not specific. The hallucinations in GPT-4 is the lowest (54.4%). Importantly, GPT-4 produced only names of fungi which actually produce chemical compounds.
@@ 5.6 Limitations
We recognise the following limitations of our study. First, LLMs are evaluated on a specific biological dataset related to chemical compounds, fungi and antibiotic activity. The models can perform differently for other biological relation. The sample size of 10 compounds is relatively small and increasing the size would lead to more representative results. However, due to a significant effort, including domain experts, required to evaluate generated texts, we argue that our findings highlight current limitations and challenges regarding LLM’s applications in the biomedical domain.
@@
Although we evaluated 19 prompt designs, the prompt engineering could be improved with e.g. a systematic search or optimal prompt algorithms. No fine-tuning, nor few-shot learning was performed. Evaluated LLM’s are of-the-shelf models. Moreover, this analysis concentrated on comparing and contrasting the behaviour of different models with regard to their factuality for a focused task within a scientific domain of discourse. Models have a diverse set of corpora, number of parameters and design intent, and it was not the intent of this study equalise and control for these settings. Instead, this analysis aimed to sample a highly diverse set of models, in order to elicit, which of these features are relevant for factual, knowledge-based inference within LLMs.
@@
A total of $>$ 150 000 fungi species have been described in the literature [42]. The naming of fungal species is subject to change over time [42, 43], and moreover, the number of fungal names continue to increase over time. Despite the improvements in information sharing, the changing fungal nomenclature rules [44, 45] and providing a tool for name standardisation, the existing literature still contains plenty of synonyms, isonyms, homonyms, orthographic variants and misused names that are not in accordance with the standard nomenclature system [46]. In addition, some taxon names in the databases may be variants, synonyms or invalidly published names for various reasons. This may lead to inaccurate search results and renders the assessment of factual knowledge in generated relations challenging even for the domain expert.
6 Related Work
Large Language models achieve state-of-the-art performance in many NLP related-task, sometimes outperforming previous baselines in few-shot settings [7]. However, the factuality of the knowledge that can be extracted by prompting these models has been widely discussed. Since the first demonstration by Petroni et al. [47] on Masked Language Models (MLM), several studies have examined the mechanisms and biases underlying these predictions, questioning their applicability beyond their potential as KBs.
Cao et al. [48] showed that predictions from MLM are highly prompt-biased and that, while external context apparently helps the models, its mainly via type guidance or answer leakage. In follow-up works, many domain-specific datasets and original prompting approaches were developed for factual knowledge extraction on MLM [49, 50, 51, 52, 53, 54, 16, 55]. In the biomedical area, Sung et al. [56] extended the work of Cao et al. on a new dataset, bioLAMA, confirming their observations for biological knowledge. More recently, Kalo and Fichtel [57] extended the evaluation on a novel benchmark dataset (KAMEL) and to generative Language Models from the GPT and OPT family. Si et al. [58] evaluated the generalisability, fairness, calibration, and factuality of different prompting strategies on GPT-3. According to Mallen et al. [59], it has been observed that retaining less popular factual knowledge is hard for LLM and simply increasing their size does not significantly improve their ability to remember. Razniewski et al. [60] presented qualitative arguments that implicit LMs are not suitable as substitutes for explicit KBs, but may play a major role in the augmentation and curation of KBs.
Howard et al. [61] examined the implications of generative AI for antimicrobial advisory in infection scenario-based questions and outlined the limitations of implementing ChatGPT for interventions due to deficits in situational awareness, inference, and consistency. Li et al. [62] reported that ChatGPT performs moderately or poorly in several biomedical tests and is unreliable in actual clinical use. Furthermore, implementing ChatGPT into clinical practice is associated with numerous obstacles, such as deficits in situational awareness, reasoning, and consistency [63]. While the use of natural language processing in healthcare is not novel, the introduction of GPT-4 has elicited intense debate regarding its potential opportunities and challenges in healthcare [64].
This paper builds upon findings from several key studies to address the limitations and potentials of LLMs in scientific knowledge encoding. Zhao et al.(2021) [65], Kassner, Krojer, and Schutze (2020) [66] and Wysocki et al. (2023) [16] emphasize the inconsistency of language models and their tendency to replicate prevalent responses, a form of co-occurrence bias also noted in LLMs by Kandpal et al. (2023) [67] and Kang and Choi (2023) [68]. This tendency undermines the models’ ability to generalize, a concern supported by Razeghi et al. (2022) [69], who report an accuracy disparity tied to the frequency of terms in training data. Biderman et al. (2023) [70], Power et al. (2022) [71] and Tirumala et al. (2022) [72] explore delicate balance between memorization’s beneficial role in preventing the hallucinations, versus outputting entire sequences from their training data verbatim, as a key safety concern. Wang et al. [73] and Delmas, Wysocka, and Freitas (2023) [74] remark on LLMs’ underperformance when compared to supervised settings, positing retrieval-augmentation as a potential remedy for knowledge-dense tasks. Together, these studies highlight the need for strategies to refine LLMs’ bias and factuality, enhancing their application in biomedical domain.
@@ 7 Conclusions
This work explores the potential of LLMs to encode domain-specific factual knowledge, with an emphasis on the biomedical field. To assess this, we introduced an evaluation framework aimed at mitigating the labor-intensive nature of text generation evaluation, promoting efficiency without sacrificing accuracy. Our framework scrutinizes fluency, prompt-alignment, semantic coherence, factuality, and specificity in the generated text. Is was applied to ten off-the-shelf models, focusing on two tasks: chemical compound definition generation and chemical-fungus relation determination via entity generation. The results reveal that, although recent advancements have enhanced fluency and semantic coherence in LLMs, a reliable performance in a zero-shot setting remains elusive. The factuality of the generated content, especially in biological contexts, is still lacking, with models showing a pronounced bias towards entities that are more prevalent in the training corpus. Our findings align with previous studies across various domains [69, 73, 65, 66, 67], confirming co-occurrence biases irrespective of model size. Notably, GPT-4, while outperforming others by demonstrating a form of epistemic caution, acknowledged its knowledge limitations, thereby preventing certain hallucinations. It recognized 70% of chemical compounds and produced less than half of the factual and specific relations to fungi. This compares to the Llama2 model, which recognized 20% of compounds and 12.7% of relations to fungi. Our study addresses the increasing demand for thorough assessments of LLMs in various fields, advancing beyond the standard benchmarks that are commonly employed. This research is among the first to investigate LLM performance within the biomedical domain, specifically probing whether these models simply recall information or truly comprehend the symbolic connections between chemicals and fungi when generating definitions.
@@
Future work. As a continuation of this study, we plan to further assess LLMs in the realm of biomedical inference, keeping pace with the rapid advancements in the field by including the latest models. We plan to extent our proposed framework addressing the need for efficient qualitative and quantitative evaluation of responses generated by LLMs. Such frameworks should go beyond the well-known benchmarks and be applicable in any specific biomedical applications, such as the one exemplified in this study (chemical-fungus relation). The promising future direction is to use prompts containing biomedical texts as contexts (abstract or full texts) combined with few-shot learning or in-context learning, as well as the integration with robust, high-recall retrieval augmentation mechanisms, e.g. [74].
Acknowledgments
The authors would like to thank the support of Joël Dumoulin, Joel Rossier and Colombine Verzat.
@@ Acknowledgments
Supplementary Information
Supplementary Tables
Table S.1: The dataset for the experiments contains ten chemical compounds (chem name): five with (antibiotic activity = 1) and five without antibiotic activity (antibiotic activity = 0). For each relation of a selected chemical compound with a fungus (fungi name) there is a number of references (nb ref) describing the relation based on PubChem.
@@
Table S.2: Examples of generated text for selected models, selected compounds as input, and selected prompts. In the ’Comment’ column there are correct answers or comments by the authors referring to semantic coherence or factuality of generated text.
@@
Table S.3: Examples of generated text for selected input: chemical compound Myriocin and a simple prompt: ’{entity 1}’ (for Galactica: What is the {entity 1}?) for all ten models analyzed in the paper. Factuality of generated text are provided for all generated responses.
@@
Galactica
Question: What is the Myriocin ? Answer: Myriocin
0
@@
Table S.4: Generated fungus name in Task 2: chemical-fungus relation determination via entity generation task for one chemical compound Alternariol for 11 prompts (P8-P18) (for Galactica model only P19 prompt). Fungus names that are generated correctly in bold.
Prompt GPT-2 GPT-3 GPT-neo Bloom BioGPT BioGPT-Large Galactica ChatGPT GPT-4 Llama 2-7B Llama 2-13B Llama 2-70B
Question: Which fungus
produces Alternariol ?
X X X X X X Alternaria X X X X X
Alternariol is produced by Alternaria alternata Alternaria alternata Phomopsis amygdali X Alternaria
@@ X X X X X X Alternaria X X X X X
Alternariol is produced by the fungus Alternaria alternata Penicillium acnes Alternaria alternata Phomopsis amygdali X Alternaria
@@ X Alternaria
Alternariol is produced by fungi, such as Alternaria Aspergillus Alternaria alternata Alternaria alternata X
@@
Alternariol is isolated from Alternaria Phomopsis amygdali X Alternaria Alternaria Alternaria
@@
Alternariol originally isolated from Alternaria alternata Aspergillus terreus Phellinus linteus X Alternaria tenuissima Alternaria alternata Alternaria alternata Alternaria alternata Alternaria
Alternariol is isolated from the fungus Alternaria alternata Alternaria alternata Phomopsis amygdali X Alternaria
@@
Alternariol is isolated from fungi Alternaria X Alternaria
@@
Alternaria alternata Phomopsis X
@@
Alternariol is a natural product found in X Alternaria Alternaria Alternaria alternata
@@
X Alternaria Alternaria Alternaria alternata Alternaria
Table S.5: Performance in the Task 2: chemical-fungus relation determination via entity generation task. Evaluation according to the framework (see Figure 3). 10 chemicals, 11 prompts (P8-P18) (for Galactica only P19 prompt), 110 outputs per model in total (for Galactica 10 outputs). Analysis of the occurrence of the most common fungus in chemical compound-fungus relation Aspergillus sin tested prompts.

### v3

@@ 5 Discussion
Our observations, initially rooted in the analysis of fungus-chemical interactions, hold relevance for a range of specific domains within the biomedical field. In this section, we detail these broadly applicable insights, directing the evaluator’s attention to essential concerns. Key observations include biases in LLM outputs towards more prevalent topics, difficulties in accurately identifying rare entities, a tendency for outputs to stray from the intended biomedical context, and significant variability in performance influenced by the design of the prompts.
@@ 5.1 LLMs’ biases towards compounds and fungi
5.1.1 Correctly generated relations are homogeneous and prevalent.
Considering the models with outputs that contain the fungus name, in majority they recognise one specific factual chemical-fungus relation, although the chemical can be produced by several fungi (RQ1). For instance, in all generated relations which were factual, for Fumitremorgin C the fungus was either Aspergillus fumigatus or Aspergillus (39 out of 39 correct fungi name, from 66 generated fungi in total). Factually generated relations for Ergosterol are with two fungi: Aspergillus (11/23) and Ganoderma lucidum (7/23). Interestingly, Aspergillus is output by BioGPT and Ganoderma lucidum by BioGPT-Large. Although trained on the same corpus, we hypothesise that the higher number of parameters in BioGPT-Large (347M vs 1.5B) led to a more specific relation (RQ2).
5.1.2 Bias towards Aspergillus - the most cited fungus in PubMed.
@@
We observe that the models are biased towards certain compounds (RQ1). For instance, BioGPT generates the fungus Aspergillus as the answer in most of the prompts (79%, 45/57), which is rarely correct (only 11 times; 19%) (Table A.7). We attribute this bias to the imbalance in the training corpus, as the Aspergillus is a ubiquitous fungus in home and hospital environments. The PubMed search outputs almost 60k references related to Aspergillus. The most cited compound from our analysis Ergosterol outputs $\sim$6000, second Aphidicolin $\sim$2500. Thus, we argue Aspergillus is over-represented in the training corpus and occurs as a statistically most probable fungus to be related to chemicals.
@@
5.1.3 Rare chemical compound confused with an overrepresented biological process.
@@
Another example, Conidiogenone was consistently not recognised by the best performing models (GPT-4, ChatGPT, BioGPT-large). We argue that the models confuse Conidiogenone with conidiogenesis or conidiation (RQ1). Conidiogenesis is a widespread morphogenetic process that filamentous fungi undertake for dispersion. Conidiation is also important for pathogenicity of phytopathogens. Understanding the cellular mechanism of conidiation is a highly relevant research topic, thus it has been extensively studied. Conidiogenone is associated with nine results in a PubMed search, conidiogenesis with 346 and conidiation with 6047. The strongest and widespread stimulus for conidiation among filamentous fungi is the exposure of hyphae to the air. Often the generated texts focused on this aspect of the process, suggesting that the model encoded the signal for the representation of the process instead of the chemical compound.
5.1.4 Output not focused on fungi despite the prompt context.
@@
In the case of Ergosterol, which is a compound widely known for its significance in the scientific community, the LLM specifically trained for the bio-domain (BioGPT-Large) outperforms the rest (RQ3). It generated seven factual Ergosterol-fungus relations from 11 prompts (Table A.8). For the other models, the relation to fungus was either minimally addressed or not considered at all, particularly for the models that have generally demonstrated superior efficacy (i.e., GPT-4 with 0/11 and ChatGPT with 4/11 factual responses). We observe that GPT-4 and ChatGPT focus on Ergosterol’s biosynthesis in fungi and relevance as an antifungal target, rather than on the producing fungi, despite the prompt that explicitly defines such context. Importantly, these two models do not generate text as a continuation of the prompt, but they generate new paragraphs.
@@ 5.2 Limited ability of LLMs as Knowledge Bases
In this study, we provide evidence of the limitations in the determination of factual knowledge from LLMs in the text generation tasks. Best models were able to provide only 6/10 of the factual description of chemical entities, and only 5/10 of the factual description of chemical-fungus relations. Despite small sample size in our evaluation, it clearly shows that the performance is insufficient for a systematic and reliable application (RQ1, RQ4).
@@ 5.3 High dependency and variability due to prompt design choices
A desirable property of a LLM is to a consistent answer regardless of the prompts design. For instance, the prompt ‘{entity 1} is produced by fungi, such as’ should be equivalent to ‘{entity 1} is isolated from fungi, such as’ because it refers to the same biological relation. However, we observe a variation in the output as the evidence of sensitivity for prompt design (RQ1). Similarly to Masked Language Models Zhou et al. (2023), in order to fully exploit GPTs capabilities, prompt engineering is required. In our study, we observe that providing context in the prompt improves the performance. Of note, ChatGPT and GPT-4 are the least sensitive and produce new paragraphs instead of finishing the prompt (RQ3).
@@ 5.4 Llama 2, despite large size, responses lack factuality and specificity
GPT-4 outperforms ChatGPT (as well the rest tested models) in terms of factuality of the answer (RQ3). It scores 70% (7/10) and 43.6% (48/110) for chemical entity recognition and chemical-fungi relation recognition, respectively. In general, the factual knowledge of the entire generated text was higher (still significantly higher in GPT-4), but the output did not contain the relation that was prompted for, so they were considered not specific. The hallucinations in GPT-4 are the lowest (54.4%). Importantly, GPT-4 produced only names of fungi which actually produce chemical compounds.
@@ 5.6 Limitations
We recognise the following limitations of this investigation. First, LLMs are evaluated on a specific biological dataset related to chemical compounds, fungi and antibiotic activity. The models can perform differently for other biological relations. The sample size of 10 compounds is relatively small and increasing the size would lead to more representative results.
@@
Although we evaluated 15 prompt designs, the prompt engineering could be improved with e.g. a systematic search or optimal prompt algorithms. No fine-tuning, nor few-shot learning was performed. The analysis used of-the-shelf LLMs, focusing on their factuality in a specific scientific task. Despite the models’ varying corpora, parameters, and design intents, the study did not aim to equalise these factors but rather to explore which features contribute to factual knowledge-based inference in LLMs.
@@
A total of $>$ 150 000 fungi species have been described in the literature Wang et al. (2022). The naming of fungal species is subject to change over time Wang et al. (2022); Richards et al. (2017), and moreover, the number of fungal names continue to increase over time. Despite the improvements in information sharing, the changing fungal nomenclature rules Aime et al. (2021); Turland et al. (2018) and providing a tool for name standardisation, the existing literature still contains plenty of synonyms, homonyms, orthographic variants and misused names that are not in accordance with the standard nomenclature system Lücking et al. (2020). In addition, some taxon names in the databases may be variants, synonyms or invalidly published names for various reasons. This may lead to inaccurate search results and renders the assessment of factual knowledge in generated relations challenging even for a domain expert.
@@ 6 Conclusions
This work explores the potential of LLMs to encode domain-specific factual knowledge, with an emphasis on the biomedical field. To assess this, we introduced a framework to streamline the labor-intensive process of human expert evaluation, using non-experts for preliminary assessments to optimise the time and expertise of specialists, without sacrificing accuracy and efficiency. In our study, this approach resulted in a 33% and 46% reduction in experts’ effort for two different tasks. The framework replicates the process of human evaluation, which remains an indispensable step regardless of the presence of automated evaluation systems Bavaresco et al. (2024). This is due to the fact that every user in the biomedical field will invariably wish to conduct manual checks of the model using a variety of familiar examples. Essentially, this framework serves as a directive on the aspects to focus on.
@@
The results reveal that, although recent advancements have enhanced fluency and semantic coherence in LLMs, a reliable performance of the models to operate as a domain-specific knowledge base remains elusive. The factuality of the generated content, especially in biological contexts, is still lacking, with models showing a pronounced bias towards entities that are more prevalent in the training corpus. Our findings align with previous studies across various domains Razeghi et al. (2022); Wang et al. ; Zhao et al. (2021); Kassner et al. (2020); Kandpal et al. (2023), confirming co-occurrence biases irrespective of model size. Notably, GPT-4, while outperforming others by demonstrating mechanisms of epistemic awareness, acknowledging its knowledge limitations, thereby preventing certain hallucinations. It recognised 70% of chemical compounds and produced less than half of the factual and specific relations to fungi. This compares to the Llama 2 model, which recognized 20% of compounds and 12.7% of relations to fungi. Our study addresses the increasing demand for thorough assessments of LLMs in various fields, advancing beyond the standard benchmarks that are commonly employed. This research complements existing works investigating LLM performance within the biomedical domain, specifically probing whether these models simply recall information or truly understand the symbolic connections between chemicals and fungi when generating definitions.
@@
Future work. As a continuation of this study, we plan to further assess LLMs for multi-hop biomedical inference. We plan to extend our proposed framework addressing the need for efficient qualitative and quantitative evaluation of responses generated by LLMs. Such frameworks should go beyond the well-known benchmarks and be transferable to other biomedical domains, such as the one exemplified in this study (chemical-fungus relation). A promising future direction is the use of robust retrieval augmented generation (RAG) mechanisms integration with robust, high-recall retrieval augmentation mechanisms, e.g. Delmas et al. (2023); Wysocki et al. (2024) in coordination with symbolic reasoning methods.
CRediT authorship contribution statement
Magdalena Wysocka: Writing – review and editing, Writing – original draft, Software, Methodology, Formal analysis, Data curation, Visualization, Conceptualization. Oskar Wysocki: Supervision, Writing – review and editing, Formal analysis, Methodology, Conceptualization. Maxime Delmas: Writing – review and editing, Writing – original draft, Software. Vincent Mutel: Project administration, Methodology. André Freitas: Writing – review and editing, Supervision, Investigation.
@@
Acknowledgments
J. Dumoulin, J. Rossier and C. Verzat for their support.
@@ Acknowledgments
Appendix A Supplementary data
The following is the Supplementary material related to this article.
Figure A.1: The number of outputs evaluated at each STEP of the framework in Task 1 for simple prompts, showing the percentage of results evaluated by non-expert and expert. Results are aggregated for prompts P1-P2 from the given model.
Figure A.2: The number of outputs evaluated at each STEP of the framework in Task 1 for context-based prompts, showing the percentage of results evaluated by non-expert and expert. Results are aggregated for prompts P3-P4 from the given model.
Figure A.3: The number of outputs evaluated at each STEP of the framework in Task 2, showing the percentage of results evaluated by non-expert and expert. Results are aggregated for all prompts P5-P15 from the given model.
(a) A)
(b) B)
Figure A.4: Heatmap showing the factuality (STEP 2) (A) and specificity (STEP 3) (B) of outputs across different models and prompts in Task 1. The x-axis represents the prompts P1-P4, while the y-axis lists the AI models used in the study. The intensity of the color indicates the effectiveness of a given prompt, with darker shades representing responses with higher factuality (A) or factual responses with higher specificity (B).
(a) A)
(b) B)
Figure A.5: Heatmap showing the factuality (STEP 2B) (A) and specificity (STEP 3) (B) of outputs across different models and prompts in Task 2. The x-axis represents the prompts P5-P15, while the y-axis lists the AI models used in the study. The intensity of the color indicates the effectiveness of a given prompt, with darker shades representing responses with higher factuality (A) or factual responses with higher specificity (B).
Table A.1: Table A.1. The dataset for the experiments contains ten chemical compounds (chem name): five with (antibiotic activity = 1) and five without antibiotic activity (antibiotic activity = 0). For each relation of a selected chemical compound with a fungus (fungi name) there is a number of references (nb ref) describing the relation based on PubChem.
@@ Appendix A Supplementary data
Table A.2: Table A.2. Summary of the evaluated Large Language Models ordered by the release date.
Model Lab Training Corpus Vocabulary size Model size (parameters) used/available Layers (L), hidden size (H), heads (A) Maximum input sequence length (tokens) Date of the model release* Reference
GPT-2 OpenAI WebText: 40 GB of text, 8M documents, from 45M webpages upvoted on Reddit (all Wikipedia pages removed) 50,257 124M/1.5B L=24, H=1024, A=16 1024 Feb 2019 Radford et al. [2019]
GPT-3 OpenAI 570 GB plaintext, 0.4 trillion tokens. (410B CommonCrawl, 19B WebText2, 3B English Wikipedia, and two books corpora (12B Books1 and 55B Books2)) 175B L=96, H=12288, A=96 2048 June 2020 Brown et al. [2020]
GPT-neo EleutherAI The Pile - a large scale curated dataset created by EleutherAI 825GB, 380 billion tokens 50,257 1.3B/2.7B L=24, H=2048, A=16 2048 March 2021 Gao et al. [2020]
Bloom BigScience 45 natural languages 12 programming languages In 1.5TB of pre-processed text, converted into 350B unique tokens 250,680 560M-3B/176B L=24, H=1024, A=16 2048 July 2022 BigScience [2022]
BioGPT Microsoft 15M PubMed abstracts (trained from scratch) 42,384 347M/347M L=24, H=1024, A=16 1024 Oct 2022 Luo et al. [2022]
BioGPT-large Microsoft 15M PubMed abstracts (trained from scratch) 42,384 1.5B/1.5B L=24, H=1024, A=16 2048 Oct 2022 Luo et al. [2022]
ChatGPT OpenAI unavailable 20B (175B) 4096 Nov 2022
GPT-4 OpenAI unavailable L=48, H=12288, A=96 8192 Feb 2023
7B L=32, H=4096, A=32
Llama 2 Meta AI unavailable 13B L=40, H=5120, A=40 4096 July 2023 Touvron et al. [2023]
70B L=80, H=8192, A=64
Table A.3: Table A.3. Examples of generated text for selected models, selected compounds as input, and selected prompts. In the ’Comment’ column there are correct answers or comments by the authors referring to semantic coherence or factuality of generated text.
@@
Table A.4: Table A.4. Examples of generated text for selected input: chemical compound Myriocin and a simple prompt: ’{entity 1}’ for all eleven models analyzed in the paper. Factuality of generated text are provided for all generated responses.
@@
Table A.5: Table A.5. Generated fungus name in Task 2: chemical-fungus relation determination via entity generation task for one chemical compound Alternariol for 11 prompts (P5-P15). Fungus names that are generated correctly in bold.
Prompt GPT-2 GPT-3 GPT-neo Bloom BioGPT BioGPT-Large ChatGPT GPT-4 Llama 2-7B Llama 2-13B Llama 2-70B
Alternariol is produced by Alternaria alternata Alternaria alternata Phomopsis amygdali Alternaria
@@
Alternariol is produced by the fungus Alternaria alternata Penicillium acnes Alternaria alternata Phomopsis amygdali Alternaria
@@
Alternariol is produced by fungi, such as Alternaria Aspergillus Alternaria alternata Alternaria alternata
@@
Alternariol is isolated from Alternaria Phomopsis amygdali Alternaria Alternaria Alternaria
@@
Alternariol originally isolated from Alternaria alternata Aspergillus terreus Phellinus linteus Alternaria tenuissima Alternaria alternata Alternaria alternata Alternaria alternata Alternaria
Alternariol is isolated from the fungus Alternaria alternata Alternaria alternata Phomopsis amygdali Alternaria
@@
Alternariol is isolated from fungi Alternaria Alternaria
@@
Alternaria alternata Phomopsis
@@
Alternariol is a natural product found in Alternaria Alternaria Alternaria alternata
@@
Alternaria Alternaria Alternaria alternata Alternaria
Table A.6: Table A.6. Generated fungus name in Task 2: chemical-fungus relation determination via entity generation task for one chemical compound Verrucarin A for 11 prompts (P5-P15). Fungus names that are generated correctly in bold.
Prompt GPT-2 GPT-3 GPT-neo Bloom BioGPT BioGPT-Large ChatGPT GPT-4 Llama 2-7B Llama 2-13B Llama 2-70B
Verrucarin A is produced by Verrucaria nigrescens Aspergillus niger Penicillium verrucosum Myrothecium verrucaria Myrothecium verrucaria
Penicillium verrucosum,
Aspergillus niger
Verrucaria
Verrucarin A is produced by the fungus Verrucaria nigrescens Penicillium expansum Aspergillus niger Penicillium verrucosum Myrothecium verrucaria Myrothecium verrucaria
Verrucarin A is produced by fungi Verrucaria Neurospora crassa Penicillium Myrothecium verrucaria Myrothecium verrucaria Verrucaria Verrucaria Verrucaria
Verrucarin A is produced by fungi, such as Verrucaria nigrescens
Penicillium roqueforti,
Penicillium verrucarinum
Aspergillus terreus Penicillium verrucosum
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium verrucaria Penicillium verrucosum Aspergillus
Verrucarin A is isolated from Verrucaria maura Penicillium verrucosum Myrothecium verrucaria Myrothecium verrucaria Verrucaria Verrucaria Aspergillus
Verrucarin A originally isolated from Verrucaria nigrescens Penicillium verrucosum Myrothecium verrucaria Myrothecium verrucaria Penicillium verrucarum
Verrucarin A is isolated from the fungus Verrucaria nigrescens Penicillium expansum Aspergillus versicolor Penicillium verrucosum Myrothecium verrucaria Myrothecium verrucaria
Verrucarin A is isolated from fungi Verrucaria Aspergillus Myrothecium verrucaria Myrothecium verrucaria Verrucaria Verrucaria Verrucaria
Verrucarin A is isolated from fungi, such as Verrucaria maura
Conchocalyx irregularis,
Aspergillus ochraceus,
Stachybotrys chartarum
Aspergillus terreus Penicillium verrucosum
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium verrucaria Verrucaria nigrescens
Verrucarin A is a natural product found in Verrucaria Penicillium verrucosum
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium Verrucaria
Verrucarin A is a natural chemical compound found in Verrucaria Penicillium verrucosum
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium Verrucaria Verrucaria
Table A.7: Table A.7. Performance in the Task 2: chemical-fungus relation determination via entity generation task. Evaluation according to the framework (Fig. 1). 10 chemicals, 11 prompts (P5-P15), 110 outputs per model in total. Analysis of the occurrence of the most common fungus in chemical compound-fungus relation Aspergillus in tested prompts.